# Modeling isotropic polyconvex hyperelasticity by neural networks – sufficient and necessary criteria for compressible and incompressible materials

### 28 Mar 2026 v1

@@ 7.3.2 Compressible Mielke energy
This section investigates the performance of the compressible CSSV-NN, reduced CSSV-NN and Ball NN on the energy proposed by Mielke [31] in the compressible case. Once again, $a$ was set to 10. The UInvar NN was not investigated here because this framework already showed poor performance for the previous, incompressible energies. The dataset was generated by sampling diagonal deformation gradients of the form
$\boldsymbol{F}=\textbf{diag}(\lambda_{1},\lambda_{2},\lambda_{3}),$ (58)
where each principal stretch $\lambda_{i}$ was selected from the uniformly discretized range $[0.6:0.2:2.0]$. To avoid redundant permutations, only unique stretch combinations satisfying $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ were considered, resulting in 120 distinct deformation states. The training procedure was the same as for the classical models except that the number of random weight initializations per architecture was increased to 60 and the number of maximum iterations was set to 10000 since more instances showed a poor performance and slower convergence compared to all other energies presented in [16] and this work. For this reason, all models were fine-tuned by SciPy’s SLSQP method after training with the L-BFGS-B.
The best CSSV-NN achieved a mean squared error of 1.67e-04, while the best reduced CSSV-NN only achieved a mean squared error of 2.80e-03 and the best Ball NN only 1.39e-02. Most of the stress predictions made by the Ball NN are more than 10% off from the respective training data reference and also many predictions from the reduced CSSV-NN are. Even the CSSV-NN shows some predictions that are more than 10% off, see Fig. 10 (gray areas). Nevertheless, the stress over stretch behavior – here $P_{11}$ (Fig. 11) and $P_{22}$ (Fig. 12) over $F_{11}$ for various, but fixed $F_{22}$ and $F_{33}$ combinations – is captured really well by the CSSV-NN. The larger errors occur at the boundary of the training data regime (Fig. 12 bottom middle for $F_{11}$=0.6) or as a smooth approximation of the rapid stress change of the reference model (Fig. 12 top middle). The $P_{11}$ stresses are captured perfectly by the CSSV-NN (Fig. 12). The reduced CSSV-NN and the Ball NN show some larger deviations, especially for the $P_{22}$ stresses (Fig. 12). Overall, it seems like the optimizer has problems to find the optimal solution. The original energy (Eq. (55)) can be expressed in terms of ReLu functions and thus by a six layer CSSV-NN with ReLu activation. To achieve this, six linear input combinations are defined first
@@
Motivated by this, various deeper (up to ten layers) and broader (up to 500 neurons per layer) networks were additionally investigated but did not lead to better results. To get an even better distinction of the expressivity of the three neural network models in the compressible case, one final example is conducted in the following section.
@@ 7.3.3 Compressible additive Mielke-type energy
A very similar energy to the one of Mielke is obtained, if the sum of all arguments instead of the max is taken:
@@
These observations once again underline the limited expressiveness of the Ball NN and that Ball’s polyconvexity criterion is not necessary in the compressible case. Furthermore, at least the results for the Mielke energy indicate that the reduced criterion based on the elementary symmetric polynomials of the signed singular values may also be not necessary and thus that the ?-marked relation in Fig. 2 is not valid in the compressible case.
@@ 8 Conclusion
In contrast, neural networks based on more restrictive sufficient criteria show clear limitations. The inc-/Ball NN fits classical and experimental data well but fails to approximate the Mielke energy in the compressible as well as incompressible case, substantiating that Ball’s criterion is sufficient but not necessary and can significantly restrict expressiveness. In addition, the inc-/Ball NN is computationally more expensive compared to the reduced inc-/CSSV-NN. The inc-UInvar NN performs considerably worse in all settings and does not capture essential characteristics of the underlying energies. Therefore, neither the inc-/Ball NN nor the inc-/UInvar NN should be used when accuracy and flexibility are required.
Altogether, the results demonstrate that neural network models relying on necessary and sufficient criteria for polyconvexity offer clear advantages over classical invariant-based or overly restrictive approaches. Tab. 9 provides an overview on the performance of all neural network variants for all investigated incompressible material models and Treloar’s experimental data. The inc-/CSSV-NN provides the most accurate, expressive and practically applicable framework and is therefore recommended for modeling frame-indifferent, isotropic polyconvex hyperelasticity of compressible and incompressible materials. If high efficiency is required, the reduced inc-CSSV-NN may also be suitable for incompressible materials.
@@ Appendix A Proof of Theorem 3: Sufficient polyconvexity criterion based on principal stretches
Since $\widetilde{\Psi}$ is $\Pi_{3}$-invariant, it can be identified with an isotropic and frame-indifferent function. By Eq. (65) and (28), this function can be identified with $\Psi$. Thus, Eq. (18) from Theorem 1 and obtain the polyconvexity of $\Psi$. ∎

### 13 Apr 2026 v2

@@ 7.3.2 Compressible Mielke energy
This section particularly aims to compare the performance of the compressible CSSV-NN, the reduced CSSV-NN and the Ball NN using the smooth approximation of the compressible energy proposed by Mielke [31]. The UInvar NN was not investigated here, as this framework already showed poor performance for the previous incompressible energies. Notably, the original energy (54) can be expressed in terms of ReLu functions and thus by a six layer CSSV-NN with ReLu activation. To achieve this, six linear input combinations are defined first
@@
Motivated by this, various deep (up to ten layers) and wide (up to 500 neurons per layer) ReLu networks with random weight initialization were trained on strain–energy tuples generated from the original energy (54) as an additional study. Surprisingly, the optimizer consistently failed to recover the analytical solution. Accordingly, even if the space spanned by the neural network includes the analytical solution, the applied optimization algorithm might not find it. This observation is also relevant for interpreting the results of the neural networks trained on the smooth approximation (55) presented below.
As in the previous section, $a$ was set to 10 for the smooth approximation. The dataset was generated by sampling diagonal deformation gradients of the form
$\boldsymbol{F}=\textbf{diag}(\lambda_{1},\lambda_{2},\lambda_{3}),$ (58)
where each principal stretch $\lambda_{i}$ was selected from the uniformly discretized range $[0.6:0.2:2.0]$. To avoid redundant permutations, only unique stretch combinations satisfying $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ were considered, resulting in 120 distinct deformation states. The training procedure was the same as for the classical models except that the number of random weight initializations per architecture was increased to 60 and the number of maximum iterations was set to 10000 since more instances showed a poor performance and slower convergence compared to all other energies presented in [16] and this work. For this reason, all models were fine-tuned by SciPy’s SLSQP method after training with the L-BFGS-B. In addition to the eight networks sizes presented in Section 6.1 and motivated by the analytical construction above, various deeper (up to ten layers) and wider (up to 500 neurons per layer) networks were trained with the same settings but did not lead to better results.
The best CSSV-NN achieved a mean squared error of 1.67e-04, while the best reduced CSSV-NN led to a mean squared error of 2.80e-03 and the best Ball NN to 1.39e-02. Most of the stress predictions made by the Ball NN are more than 10% off from the respective training data reference and so are many reduced CSSV-NN predictions. Even the CSSV-NN shows some predictions that are more than 10% off, see Fig. 10 (gray areas). Nevertheless, the stress over stretch behavior – here $P_{11}$ (Fig. 11) and $P_{22}$ (Fig. 12) over $F_{11}$ for various, but fixed $F_{22}$ and $F_{33}$ combinations – is captured really well by the CSSV-NN. The larger errors occur at the boundary of the training data regime (Fig. 12 bottom middle for $F_{11}$=0.6) or as a smooth approximation of the rapid stress change of the reference model (Fig. 12 top middle). The $P_{11}$ stresses are captured perfectly by the CSSV-NN (Fig. 12). The reduced CSSV-NN and the Ball NN show some larger deviations, especially for the $P_{22}$ stresses (Fig. 12). Overall, the remaining error of the CSSV-NN is most likely attributable to the optimizer converging to local minima, since the smooth target energy (55) can, in principle, be approximated arbitrarily well by the chosen network architecture and the original energy itself admits an exact representation by a CSSV-NN.
To get an even better distinction of the expressivity of the three neural network models in the compressible case, one final example is conducted in the following section.
@@ 7.3.3 Compressible additive Mielke-type energy
A very similar energy to the one of Mielke is obtained, if the maximum is replaced by the sum of all arguments:
@@
These observations once again underline the limited expressiveness of the Ball NN and that the improved version of Ball’s polyconvexity criterion is not necessary in the compressible case. Furthermore, at least the results for the Mielke energy indicate that the reduced criterion based on the elementary symmetric polynomials of the signed singular values may also be not necessary and thus the ?-marked relation in Fig. 2 seems not to be valid in the compressible case. However, a rigorous mathematical proof is still missing.
@@ 8 Conclusion
In contrast, neural networks based on more restrictive sufficient criteria show clear limitations. The inc-/Ball NN fits classical and experimental data well but fails to approximate the Mielke energy in the compressible as well as incompressible case, substantiating that both – Ball’s original criterion and the improved version – are sufficient but not necessary and can significantly restrict expressiveness. In addition, the inc-/Ball NN is computationally more expensive compared to the reduced inc-/CSSV-NN. The inc-UInvar NN performs considerably worse in all settings and does not capture essential characteristics of the underlying energies. Therefore, neither the inc-/Ball NN nor the inc-/UInvar NN should be used when accuracy and flexibility are required.
Altogether, the results demonstrate that neural network models relying on necessary and sufficient criteria for polyconvexity offer clear advantages over classical invariant-based or overly restrictive approaches. Tab. 9 summarizes the performance of all neural network variants for all investigated material models and Treloar’s experimental data. The inc-/CSSV-NN provides the most accurate, expressive and practically applicable framework and is therefore recommended for modeling frame-indifferent, isotropic polyconvex hyperelasticity of compressible and incompressible materials. If high efficiency is required, the reduced inc-CSSV-NN may also be suitable for incompressible materials.
@@ Appendix A Proof of Theorem 3: Sufficient polyconvexity criterion based on principal stretches
Since $\widetilde{\Psi}$ is $\Pi_{3}$-invariant, it can be identified with an isotropic and frame-indifferent function. By Eq. (64) and (27), this function can be identified with $\Psi$. Thus, Eq. (17) from Theorem 1 holds and one obtains the polyconvexity of $\Psi$. ∎